Let Z be a standard normal random variable. Use the calculator provided, or this table, to determine the value of c.

P≤c≤Z0.950.8154

Carry your intermediate computations to at least four decimal places. Round your answer to two decimal places.
Answer #1

P(c≤Z≤0.95)= 0.8154

Or, P( -c≤Z≤0)+P(0≤Z≤0.95)= 0.8154

Or, P(-c≤Z≤0) + 0.3289=0.8154

Or, P(-c≤Z≤0)= 0.8154-0.3289=0.4865

Or, P(0≤Z≤c) = 0.4865 [using symmmetricity , P(-c≤Z≤0)=P(0≤Z≤c)]

Or, c= 2.21 ,[ As from normal table P( 0≤Z≤2.21)=0.4865]

९६५ 2 20.१४ -24 6.च 60.5- - 0.8154 (2.4 2. 24 ८८ < 0 पुर) so C, -224